The resistance of a resistance thermometer at the ice point is 40 Ω. Its resistance at temperatures T1 and T2 are 48 Ω and 53.2 Ω respectively. What is the ratio of T2 and T1?


1
Expert's answer
2022-04-04T09:04:50-0400

Answer

Temperature at at T°C is given

"R_T=R_0(1+\alpha(T_2-T_1))"

So

"\alpha=\frac{1}{T_2-T_1}* (\frac{R_T}{R_0}-1)"

Apply in both case

"\alpha=\frac{1}{T_2-T_1}* 0.2=\frac{1}{T_2-T_1}* 0.325"

So this equation gives the ratio of temperature.